# Unified Text-Image-to-Video Generation: A Training-Free Approach to Flexible Visual Conditioning

### v2

@@ Abstract
Text-image-to-video (TI2V) generation is a critical problem for controllable video generation using both semantic and visual conditions. Most existing methods typically add visual conditions to text-to-video (T2V) foundation models by finetuning, which is costly in resources and only limited to a few pre-defined conditioning settings. To tackle these constraints, we introduce a unified formulation for TI2V generation with flexible visual conditioning. Furthermore, we propose an innovative training-free approach, dubbed FlexTI2V, that can condition T2V foundation models on an arbitrary amount of images at arbitrary positions. Specifically, we firstly invert the condition images to noisy representation in a latent space. Then, in the denoising process of T2V models, our method uses a novel random patch swapping strategy to incorporate visual features into video representations through local image patches. To balance creativity and fidelity, we use a dynamic control mechanism to adjust the strength of visual conditioning to each video frame. Extensive experiments validate that our method surpasses previous training-free image conditioning methods by a notable margin. Our method can also generalize to both UNet-based and transformer-based architectures.
@@ 1 Introduction
  • We introduce FlexTI2V, a novel training-free approach to incorporate flexible image conditioning into pretrained text-to-video diffusion models. The key innovation is a random patch swapping method to incorporate visual guidance into the denoising process of video generation in a plug-and-play manner.
  • The experiments indicate that our approach notably surpasses the existing training-free image conditioning methods. We further validate the generalization of our method to both UNet and transformer architectures.
@@ 2 Related Work
Video Generation Conditioned on Texts and Images Text-to-video (T2V) generation has gained impressive advancement since the emergence of diffusion models [ho2022video, he2022latent, hong2022cogvideo, blattmann2023align, kong2024hunyuanvideo, yang2025cogvideox, wang2025wan, wang2023modelscope]. Given the ambiguity of language in describing visual details, RGB images are exploited as the condition for video generation. Image conditioning can be used alone for image-to-video (I2V) generation [blattmann2021understanding, dorkenwald2021stochastic, wang2022latent, mahapatra2022controllable, blattmann2023stable, li2024generative], or integrated with textual prompts for text-image-to-video (TI2V) generation [ni2023conditional, zhang2023i2vgen, jiang2024videobooth, chen2024livephoto, wang2025dreamvideo, huang2025step]. Here, we mainly discuss previous TI2V methods. hu2022make propose TI2V task for the first time, and develop a transformer-based architecture to achieve decent controllability and diversity in generated videos. guo2023animatediff develop a practical framework, AnimateDiff, to train a model-agnostic plug-and-play motion module that can turn any personalized text-to-image diffusion model into an image animator. xing2024dynamicrafter introduce a novel model – DynamiCrafter, which can animate both human bodies and objects. fu2023tell propose a unified video completion method that can tackle image animation, rewinding and infilling using a single model. Most prior methods require training some layers or the entire model to incorporate image conditioning or motion over time. As a result, existing models can only handle TI2V tasks pre-defined in training, such as animation and infilling. In this paper, we introduce a generic TI2V problem that unifies previous TI2V tasks. Our proposed method is able to condition T2V diffusion models on an arbitrary number of images at arbitrary positions of the output videos without finetuning.
@@
Video Frame Interpolation Video frame interpolation has been extensively studied in recent years, with the model architectures spanning across convolutional networks [niklaus2017video, liu2020enhanced, reda2022film], transformers [fu2023tell] and diffusion models [feng2024explorative, github2024cogvideox, github2025wan]. liu2017video propose a convolutional network to synthesize video frames by flowing pixel values from bounded images. xu2019quadratic point out that most interpolation methods assume a uniform motion between two consecutive images, thus resulting in oversimplified linear models for interpolation. To generate complex motion flows, They propose a quadratic interpolation method to model acceleration information in videos. Previous work mainly aims at finding a plausible path of transition from the first video frame to the last one. How to go beyond the two bounded images remains understudied. In this work, we work on the generic TI2V problem including both classic video frame interpolation and frame synthesis beyond the bound of start and end images.
@@
Training-free Methods for Diffusion Foundation Models The high computing overhead and large amounts of training data of finetune the foundation models are inevitable bottlenecks for downstream tasks. Recently, the development of training-free approaches provide an excellent trade-off between performance and cost [hertz2022prompt, mokady2023null, tumanyan2023plug, liu2024towards, khachatryan2023text2video, kara2024rave, ni2024ti2v, titov2024guide, feng2024explorative, zhou2024storydiffusion, hou2025training, liang2025looking, liang2025movie, hsiao2025tf]. hertz2022prompt introduce prompt-to-prompt method to edit synthetic images by manipulating self-attention maps, which is further extended to real photo editing using null-text inversion [mokady2023null]. kara2024rave adapt a pretrained text-guided image editing model for video editing tasks by random noise shuffling. In terms of image-conditioned video generation, the work most related to our method is TI2V-Zero [ni2024ti2v]. They duplicate the condition image and synthesize only one video frame in a single denoising process, which enforces the pretrained temporal layer to leverage visual features from the image. However, this method suffers from slow inference speed and compounding error. feng2024explorative convert an image-to-video model to a bounded-image interpolation model by merging the two separate denoising paths conditioned on the bounded images. hou2025training propose to map a single image to 3D point cloud and then synthesize new videos from other viewpoints. Existing training-free TI2V methods usually focus on a specific problem setting. In this work, we make the first attempt to develop a training-free method for various image conditioning settings, with high flexibility in image amounts and positions.
@@ 3 Preliminaries
Figure 3: Overview of the proposed FlexTI2V approach. We invert the condition image embedding to noisy representation $\tilde{\bm{x}}_{t}$ at each step. The final noise $\tilde{\bm{x}}_{T}$ is reused as initialization for video synthesis. At step $t$, we directly replace the video frames with images at the desired positions. Then, for each video frame, we randomly swap a portion of patches with bounded condition images based on the relative distance between the frame and images. Though we show a special case of using two condition images in this figure, our method can naturally extend to any number of images at any positions. Note that all operations of our method occur in the latent space. We visualize RGB images and frames on the latent representations simply for intuitive understanding.
@@
$\bm{z}_{t-1}=$ $\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\left(\bm{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\bm{\varepsilon}_{\theta}(\bm{z}_{t},t)\right)$ (2)
$+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\bm{\varepsilon}_{\theta}(\bm{z}_{t},t)+\sigma_{t}\cdot\bm{\epsilon}^{\prime}$
@@ 4 FlexTI2V
The setting of the proposed generic text-image-to-video (TI2V) generation problem is illustrated in Fig. 2. The input includes a textual prompt $y$, NN images $\bm{x}=\{x^{(n)}\}_{n=0}^{N-1}$ as visual conditions, and the expected position index $\bm{p}=\{p^{(n)}\}_{n=0}^{N-1}$ of each image in the output video. The goal is to synthesize a video consisting of $M$ frames $\bm{v}=\{v^{(m)}\}_{m=0}^{M-1}$ with each condition image at the expected position (i.e., $v^{(p^{(n)})}=x^{(n)}$ where $p^{(n)}\in[0,M-1]$). The output video is supposed to show a reasonable transition across condition images following the textual prompt. Our problem setting unifies many existing TI2V tasks in one formulation, such as image animation ($N=1$, $\bm{p}=\{0\}$), video rewinding ($N=1$, $\bm{p}=\{M-1\}$) and video frame interpolation ($N=2$, $\bm{p}=\{0,M-1\}$).
The major challenge of training-free TI2V generation is harmonizing the distinct denoising directions controlled by texts and images solely relying on frozen model parameters. The experiments of MAE [he2022masked] validate that global image information can be carried by and inferred from a portion of local patches. Inspired by this finding, the key insight of our FlexTI2V method is contextualizing each synthetic video frame with visual conditioning, by randomly swapping a portion of patches between video frames and condition images in the latent space at early denoising steps. To balance the fidelity to condition images and creativity in motion generation, we use a dynamic control strategy for the percentage of swapped patches and denoising steps for each video frame, based on their distance to the condition images. This strategy also introduces position information of condition images into the denoising process, thus making the diffusion model capable of handling multiple condition images at any desired positions.
@@ 4.1 Condition Image Inversion and Replacement
Image Inversion The overview of our method is illustrated in Fig. 3. First of all, we encode condition images $\bm{x}$ into latent embedding $\tilde{\bm{x}}$ using the pretrained encoder $\mathcal{E}$, e.g., $\tilde{\bm{x}}=\mathcal{E}(\bm{x})=\{\tilde{x}^{(n)}\}_{n=0}^{N-1}$. To fuse the visual condition with T2V denoising process, we invert the clean latent image embedding into noisy representations at each denoising step. We denote the representation at step $t=0$ as $\tilde{\bm{x}}_{0}=\tilde{\bm{x}}$ for consistent notations in diffusion models. The widely-used DDIM inversion [mokady2023null, titov2024guide] is not applicable to our problem because the video diffusion model can not estimate the noise accurately in a single image without temporal information. Instead, we use DDPM-based inversion as an alternative following [meng2022sdedit, ni2024ti2v], which directly adds Gaussian noise to $\tilde{\bm{x}}_{0}$ like diffusion process, $\tilde{\bm{x}}_{t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\;\tilde{\bm{x}}_{0},\ (1-\bar{\alpha}_{t})\bm{I}\right)$. In denoising process, T2V diffusion models typically initialize the noise by drawing a random sample from Gaussian distribution [ho2020denoising]. Recent work has validated reusing noise inverted from images preserves the global structure in image editing [tumanyan2023plug, mokady2023null, liu2024towards], and leads to consistent video frames in video generation [ni2024ti2v]. Hence, we duplicate the inverted noise $\tilde{\bm{x}}_{T}$ at the final step, and stack them along the temporal dimension as the initial $\bm{z}_{T}$ for video denoising.
@@
$\verb+FrameReplace+\left(\tilde{\bm{x}}_{t},\bm{z}_{t},\bm{p}\right)=$ (3)
$\left\{z_{t}^{(0)},z_{t}^{(1)},\cdots,z_{t}^{(p^{(n)}-1)},\tilde{x}_{t}^{(p^{(n)})},z_{t}^{(p^{(n)}+1)},\cdots,z_{t}^{(M-2)},z_{t}^{(M-1)}\right\},$
@@ 4.2 Random Patch Swapping
Methods Image Animation Methods Video Frame Interpolation
[…]
DC [xing2024dynamicrafter] 26.20 210.77 832.38 29.86 89.10 80.95 DC [xing2024dynamicrafter] 16.20 310.74 731.90 29.69 84.26 78.68
TI2V-Zero [ni2024ti2v] 28.40 189.49 586.54 29.84 88.01 82.29 TRF [feng2024explorative] 32.20 247.88 631.54 29.47 91.99 83.35
FlexTI2V 45.40 125.49 558.29 30.97 94.40 84.38 FlexTI2V 51.60 98.41 406.30 30.92 92.34 86.88
Methods Outpainting Methods Rewinding
[…]
TI2V-Zero [ni2024ti2v] 38.60 216.80 641.07 28.80 89.01 85.01 TI2V-Zero [ni2024ti2v] 39.80 199.01 611.60 29.77 89.21 81.02
FlexTI2V 61.40 167.46 632.71 30.75 95.89 88.34 FlexTI2V 60.20 134.98 596.96 30.95 94.58 85.11
Table 1: Quantitative comparison with previous methods. “DC” is short for DynamiCrafter. “User” denotes the user study results. $\downarrow$ means a lower score in this metric indicates a better performance. The orange row refers to our method.
DC [xing2024dynamicrafter] TI2V-Zero [ni2024ti2v] TRF [feng2024explorative] FlexTI2V
Inference Time (s) 9.45 275.14 60.05 2.71
Table 2: Comparison of model efficiency in inference. “DC” is short for DynamiCrafter. The orange column refers to our method.
Random patch swapping is essentially replacing the element of $m$-th video frame at position $(i,j)$ with the element of $n$-th condition image at the same position. Each element in the latent space corresponds to a patch in RGB space. Then RPS can be formulated as updating the values in $\hat{z}_{t}^{(m)}$ following this equation:
$\footnotesize\hat{z}_{t}^{(m)}[i,j]=\mathcal{M}_{t}^{(m,n)}[i,j]\cdot\tilde{x}_{t}^{(n)}[i,j]+\left(1-\mathcal{M}_{t}^{(m,n)}[i,j]\right)\cdot\hat{z}_{t}^{(m)}[i,j].$ (4)
In this way, the visual cues from condition images are delivered to video frames through swapped patches, and then harmonized with video features by the pretrained diffusion foundation models. Each video frame only swaps patches with the two bounded condition images (the closest condition images before and after it). If all condition images are on the same side of a video frame, this frame only swaps patches with one condition image. We find a large percentage of swapped patches may include overly strong spatial alignment prior, while a small percentage contains insufficient visual features for effective guidance. To make a trade-off, we dynamically reduce the percentage and steps of RPS when the video frame is far from the position of condition image. The percentage $P(m,n,t)$ is written as
$P(m,n,t)=$ $\left\{\begin{array}[]{ll}P_{0}-\delta_{1}\cdot|m-p^{(n)}|,&0\leq t\leq\tilde{t}\\ 0,&\tilde{t}<t\leq T,\end{array}\right$. (5)
$\hskip 18.49988pt\mathrm{where}\qquad\tilde{t}=t_{0}-\delta_{2}\cdot|m-p^{(n)}|.$
$P_{0}$ and $t_{0}$ are initial swapping percentage and steps, and $\delta_{1}$, $\delta_{2}$ are decreasing strides. Our dynamic control strategy is based on the intuitive hypothesis that video frames around condition images are also perceptually similar, thus demanding more visual cues by patch swapping. Likewise, for frames far from condition images, we reduce constraints of visual conditioning accordingly to encourage more creativity in motion generation. This method also introduces position information of images into a T2V model, which makes our approach applicable to any number of images at any positions. The entire algorithm formulation is shown in Algorithm A in the supplementary.

### v3

@@ Abstract
Text-image-to-video (TI2V) generation is a critical problem for controllable video generation using both semantic and visual conditions. Most existing methods typically add visual conditions to text-to-video (T2V) foundation models by finetuning, which is costly in resources and only limited to a few pre-defined conditioning settings. To tackle these constraints, we introduce a unified formulation for TI2V generation with flexible visual conditioning. Furthermore, we propose an innovative training-free approach, dubbed FlexTI2V, that can condition T2V foundation models on an arbitrary amount of images at arbitrary positions. Specifically, we firstly invert the condition images to noisy representation in a latent space. Then, in the denoising process of T2V models, our method uses a novel random patch swapping strategy to incorporate visual features into video representations through local image patches. To balance creativity and fidelity, we use a dynamic control mechanism to adjust the strength of visual conditioning to each video frame. Extensive experiments validate that our method surpasses previous training-free image conditioning methods by a notable margin. Our method can also generalize to both UNet-based and transformer-based architectures across various image conditioning settings.
@@ 1 Introduction
  • $\bullet$
    We introduce FlexTI2V, a novel training-free approach to incorporate flexible image conditioning into pretrained text-to-video diffusion models. The key innovation is a random patch swapping method to incorporate visual guidance into the denoising process in a plug-and-play manner.
  • $\bullet$
    The experiments indicate that our approach notably surpasses the existing training-free image conditioning methods. Our method also generalizes seamlessly to both UNet and transformer architectures.
@@ 2 Related Work
Video Generation Conditioned on Texts and Images Text-to-video (T2V) generation has gained impressive advancement since the emergence of diffusion models [ho2022video, he2022latent, hong2022cogvideo, blattmann2023align, kong2024hunyuanvideo, yang2025cogvideox, wang2025wan, wang2023modelscope]. Given the ambiguity of language in describing visual details, RGB images are exploited as the condition for video generation. Image conditioning can be used alone for image-to-video (I2V) generation [blattmann2021understanding, dorkenwald2021stochastic, wang2022latent, mahapatra2022controllable, blattmann2023stable, li2024generative], or integrated with textual prompts for text-image-to-video (TI2V) generation [ni2023conditional, zhang2023i2vgen, jiang2024videobooth, chen2024livephoto, wang2025dreamvideo, huang2025step]. Here, we mainly discuss previous TI2V methods. Hu et al. [hu2022make] propose TI2V task for the first time, and develop a transformer-based architecture to achieve decent controllability and diversity in generated videos. Guo et al. [guo2023animatediff] develop a practical framework, AnimateDiff, to train a model-agnostic plug-and-play motion module that can turn any personalized text-to-image diffusion model into an image animator. Xing et al. [xing2024dynamicrafter] introduce a novel model – DynamiCrafter, which can animate both human bodies and objects. Fu et al. [fu2023tell] propose a unified video completion method that can tackle image animation, rewinding and infilling using a single model. Most prior methods require training some layers or the entire model to incorporate image conditioning or motion over time. As a result, existing models can only handle TI2V tasks pre-defined in training, such as animation and infilling. In this paper, we introduce a generic TI2V problem that unifies previous TI2V tasks. Our proposed method is able to condition T2V diffusion models on an arbitrary number of images at arbitrary positions of the output videos without finetuning.
@@
Video Frame Interpolation Video frame interpolation has been extensively studied in recent years, with the model architectures spanning across convolutional networks [niklaus2017video, liu2020enhanced, reda2022film], transformers [fu2023tell] and diffusion models [feng2024explorative, github2024cogvideox, github2025wan]. Liu et al. [liu2017video] propose a convolutional network to synthesize video frames by flowing pixel values from bounded images. Xu et al. [xu2019quadratic] point out that most interpolation methods assume a uniform motion between two consecutive images, thus resulting in oversimplified linear models for interpolation. To generate complex motion flows, They propose a quadratic interpolation method to model acceleration information in videos. Previous work mainly aims at finding a plausible path of transition from the first video frame to the last one. How to go beyond the two bounded images remains understudied. In this work, we work on the generic TI2V problem including both classic video frame interpolation and frame synthesis beyond the bound of start and end images.
@@
Training-free Methods for Diffusion Foundation Models The high computing overhead and large amounts of training data of finetune the foundation models are inevitable bottlenecks for downstream tasks. Recently, the development of training-free approaches provide an excellent trade-off between performance and cost [hertz2022prompt, mokady2023null, tumanyan2023plug, liu2024towards, khachatryan2023text2video, kara2024rave, ni2024ti2v, titov2024guide, feng2024explorative, zhou2024storydiffusion, hou2025training, liang2025looking, liang2025movie, hsiao2025tf]. Hertz et al. [hertz2022prompt] introduce prompt-to-prompt method to edit synthetic images by manipulating self-attention maps, which is further extended to real photo editing using null-text inversion [mokady2023null]. Kara et al. [kara2024rave] adapt a pretrained text-guided image editing model for video editing tasks by random noise shuffling. In terms of image-conditioned video generation, the work most related to our method is TI2V-Zero [ni2024ti2v]. They duplicate the condition image and synthesize only one video frame in a single denoising process, which enforces the pretrained temporal layer to leverage visual features from the image. However, this method suffers from slow inference speed and compounding error. Feng et al. [feng2024explorative] convert an image-to-video model to a bounded-image interpolation model by merging the two separate denoising paths conditioned on the bounded images. Hou et al. [hou2025training] propose to map a single image to 3D point cloud and then synthesize new videos from other viewpoints. Existing training-free TI2V methods usually focus on a specific problem setting. We make the first attempt to develop a training-free method for various image conditioning settings, with high flexibility in image amounts and positions.
@@ 3 Preliminaries
$\bm{z}_{t-1}=\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\left(\bm{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\bm{\varepsilon}_{\theta}(\bm{z}_{t},t)\right)+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\bm{\varepsilon}_{\theta}(\bm{z}_{t},t)+\sigma_{t}\cdot\bm{\epsilon}^{\prime}$ (2)
@@
Figure 3: Overview of the proposed FlexTI2V approach. We invert the condition image embedding to noisy representation $\tilde{\bm{x}}_{t}$ at each step. The final noise $\tilde{\bm{x}}_{T}$ is reused as initialization for video synthesis. At step $t$, we directly replace the video frames with images at the desired positions. Then, for each video frame, we randomly swap a portion of patches with bounded condition images based on the relative distance between the frame and images. Though we show a special case of using two condition images in this figure, our method can naturally extend to any number of images at any positions. Note that all operations of our method occur in the latent space. We visualize RGB images on the latent representations simply for intuitive understanding.
@@ 4 FlexTI2V
The setting of the proposed generic text-image-to-video (TI2V) generation problem is illustrated in Fig. 2. The input includes a textual prompt $y$, NN images $\bm{x}=\{x^{(n)}\}_{n=0}^{N-1}$ as visual conditions, and the expected position index $\bm{p}=\{p^{(n)}\}_{n=0}^{N-1}$ of each image in the output video. The goal is to synthesize a video consisting of $M$ frames $\bm{v}=\{v^{(m)}\}_{m=0}^{M-1}$ with each condition image at the expected position (i.e., $v^{\left(p^{(n)}\right)}=x^{(n)}$ where $p^{(n)}\in[0,M-1]$). The output video is supposed to show a reasonable transition across condition images following the textual prompt. Our problem setting unifies many existing TI2V tasks in one formulation, such as image animation ($N=1$, $\bm{p}=\{0\}$), video rewinding ($N=1$, $\bm{p}=\{M-1\}$) and video frame interpolation ($N=2$, $\bm{p}=\{0,M-1\}$).
The major challenge of training-free TI2V generation is harmonizing the distinct denoising directions controlled by texts and images solely relying on frozen model parameters. The experiments of MAE [he2022masked] validate that global image information can be carried by and recovered from a portion of local patches. Inspired by this finding, the key insight of our FlexTI2V method is contextualizing each synthetic video frame with visual conditioning, by randomly swapping a portion of patches between video frames and condition images in the latent space at early denoising steps. To balance the fidelity to condition images and creativity in motion generation, we use a dynamic control strategy for the percentage of swapped patches and denoising steps for each video frame, based on their distance to the condition images. This strategy also introduces position information of condition images into the denoising process, thus making the diffusion model capable of handling multiple condition images at any desired positions.
@@ 4.1 Condition Image Inversion and Replacement
Image Inversion The overview of our method is illustrated in Fig. 3. First of all, we encode condition images $\bm{x}$ into latent embedding $\tilde{\bm{x}}$ using the pretrained encoder $\mathcal{E}$, e.g., $\tilde{\bm{x}}=\mathcal{E}(\bm{x})=\{\tilde{x}^{(n)}\}_{n=0}^{N-1}$. To fuse the visual condition with T2V denoising process, we invert the clean latent image embedding into noisy representations at each denoising step. We denote the representation at step $t=0$ as $\tilde{\bm{x}}_{0}=\tilde{\bm{x}}$ for consistent notations in diffusion models. The widely-used DDIM inversion [mokady2023null, titov2024guide] is not applicable to our problem because the video diffusion model can not estimate the noise accurately in a single image without temporal information. Instead, we use the straightforward DDPM-based inversion as an alternative following [meng2022sdedit, ni2024ti2v], which directly adds Gaussian noise to $\tilde{\bm{x}}_{0}$ like diffusion process, $\tilde{\bm{x}}_{t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\;\tilde{\bm{x}}_{0},\ (1-\bar{\alpha}_{t})\bm{I}\right)$. In denoising process, T2V diffusion models typically initialize the noise by drawing a random sample from Gaussian distribution [ho2020denoising]. Recent work has validated reusing noise inverted from images preserves the global structure in image editing [tumanyan2023plug, mokady2023null, liu2024towards], and leads to consistent video frames in video generation [ni2024ti2v]. Hence, we duplicate the inverted noise $\tilde{\bm{x}}_{T}$ at the final step, and stack them along the temporal dimension as the initial $\bm{z}_{T}$ for video denoising.
@@
$\scriptsize\text{{FrameReplace}}\left(\tilde{\bm{x}}_{t},\bm{z}_{t},\bm{p}\right)=\left\{z_{t}^{(0)},z_{t}^{(1)},\cdots,z_{t}^{\left(p^{(n)}-1\right)},\tilde{x}_{t}^{\left(p^{(n)}\right)},z_{t}^{\left(p^{(n)}+1\right)},\cdots,z_{t}^{(M-2)},z_{t}^{(M-1)}\right\},$ (3)
@@ 4.2 Random Patch Swapping
Random patch swapping is essentially replacing the element of $m$-th video frame at position $(i,j)$ with the element of $n$-th condition image at the same position. Each element in the latent space corresponds to a patch in RGB space depending on the encoder compression ratio. Then RPS can be formulated as updating the values in $\hat{z}_{t}^{(m)}$ following the equation below:
$\hat{z}_{t}^{(m)}[i,j]=\mathcal{M}_{t}^{(m,n)}[i,j]\cdot\tilde{x}_{t}^{(n)}[i,j]+\left(1-\mathcal{M}_{t}^{(m,n)}[i,j]\right)\cdot\hat{z}_{t}^{(m)}[i,j].$ (4)
In this way, the visual cues from condition images are delivered to video frames through swapped patches, and then harmonized with video features by the pretrained diffusion foundation models. Each video frame only swaps patches with the two bounded condition images (the closest condition images before and after it). If all condition images are on the same side of a video frame, this frame only swaps patches with the single nearest condition image. We find that a large percentage of swapped patches may include overly strong spatial alignment prior, while a small percentage contains insufficient visual features for effective guidance. To make a trade-off, we dynamically reduce the percentage and steps of RPS when the video frame is far from the position of condition image. The percentage $P(m,n,t)$ is written as
$\small P(m,n,t)=\left\{\begin{array}[]{ll}P_{0}-\delta_{1}\cdot|m-p^{(n)}|,&T-\tilde{t}<t\leq T,\\ 0,&0\leq t\leq T-\tilde{t},\end{array}\right.\ \mathrm{where}\quad\tilde{t}=t_{0}-\delta_{2}\cdot|m-p^{(n)}|.$ (5)
$P_{0}$ and $t_{0}$ are initial swapping percentage and steps, and $\delta_{1}$, $\delta_{2}$ are decreasing strides. Our dynamic control strategy is based on the intuitive hypothesis that video frames around condition images are also perceptually similar, thus demanding more visual cues by patch swapping. Likewise, for frames far from condition images, we reduce constraints of visual conditioning accordingly to encourage more creativity in motion generation. We only swap patches in the first $\tilde{t}$ denoising steps, and let the model to harmonize and smooth artifacts in the remaining $T-\tilde{t}$ steps by setting $P(m,n,t)=0$. Additionally, This method also introduces position information of images into a T2V model, which makes our approach applicable to any number of images at any positions. The entire algorithm formulation is shown in the supplementary.